OpenStudy (anonymous):

Three friends went shopping. Together, Dan and Jan spent$30. Jan and Van spent $17. Dan and Van spent $25. How much did each friend spend? Please show work.

OpenStudy (anonymous):

d+j=30 J+v=17 d+v=25 -------- j=30-d v=25-d (30-d)+(25-d)=17 d=19 -------- 19+j=30 j=11 ---- 11+v=17 v=6